OpenStudy (anonymous):

Solve by the elimination method 4x-5y=-3 -8x+10y=6

OpenStudy (callisto):

4x-5y=-3 ---(1) -8x+10y=6 ---(2) (1)x2 + (2): 2(4x-5y) + (-8x+10y) = 2(-3) + (6) Can you solve it now?

OpenStudy (anonymous):

still not!

OpenStudy (anonymous):

I got that 4x+10y=6 24x-13y=-13

OpenStudy (callisto):

Oops! That doesn't look good at all.. But still you can try first. You'll see what happens! Method of elimination: 1. choose a variable. 2. find their LCM 3. add/subtract the equation so that one variable in the equation can be cancelled out 4. Solve the remaining variable, if you can.

OpenStudy (callisto):

How did you get :I got that 4x+10y=6 24x-13y=-13 ??

OpenStudy (anonymous):

I found and example online and I did it as the example

OpenStudy (callisto):

Does it related to the question you posted? I mean direct relations

OpenStudy (anonymous):

not really. the example was: 4x + 10y = 2, 3x + 5y = 5?

OpenStudy (callisto):

Alright. So, would you like solve the example you found first, or the question you posted here first? (PS: Sorry for my poor grammar today!)

OpenStudy (callisto):

*would you like to solve

OpenStudy (anonymous):

the question I posted first

OpenStudy (callisto):

Okay. Now you have a choice - do you prefer eliminate x or y in this question? For convenience, the equations are copied here again. 4x-5y=-3 -8x+10y=6

OpenStudy (anonymous):

let's say x

OpenStudy (callisto):

Okay. You want eliminate x, that means you want x to disappear when we combine the equations and solve it. As mentioned before, when we use method of elimination, we add/subtract equations in order to cancel a variable (well, some variables sometimes, but in this case, one variable). So, how can we cancel the variable x in this question - add the equations or subtract the equations?

OpenStudy (callisto):

Hmm.. Do you understand what I've written/asked? If you don't, do tell me!

OpenStudy (anonymous):

hey there the answer is 0 and by the way question cannot be of this type.

OpenStudy (callisto):

The answer is 0? Are you sure? I got different answer from yours! :|